x^2+x+1=0
x^2+x+5=0

Answer by stanbon(48546)

x^2+x+1=0
Use the quadratic formula to get:
x=[-1+sqrt(-4)]/2 or x= [-1-sqrt(-4)]/2
x=[-1+2i]/2 or x=[-1-2i]/2
--------------------------------------
x^2+x+5=0
Use the quadratic formula to get:
x=[-1+sqrt(-20)]/2 or x=[-1-sqrt(-20)]/2
x=[-1+2isqrt5]/2 or x=[-1-sqrt2isqrt(5)]/2
-----------------------------
Cheers,
Stan H.